# Assessing Logical Reasoning Capabilities of Encoder-Only Transformer Models

### v1

@@ 5 Cross-Probing Fine-Tuned LMs
LMs can successfully solve hypothesis classification tasks that involve logical reasoning. However, it raises the question: has this logical reasoning ability been acquired during the fine-tuning stage, or was it present from pretraining? Most importantly, have LMs truly developed a generalized logical reasoning capability? To examine this, we run a cross-probing task, where the LMs fine-tuned on our four logical datasets, as well as a pretrained LM, are probed on these same datasets.
Given the large number of possible tests, we restricted our investigation to a single model, RoBERTa-large, so as to dig deeper on it. This LM demonstrated a suitable balance between performance, consistency among datasets, and training time in the fine-tuning tests.
For the cross-probing task, we took the best fine-tuned RoBERTa-large model for each dataset and removed their classification heads, leaving just the transformer blocks, as in a pretrained model. We then trained a probe on the LMs’ final representations, while freezing the models’ bodies; we also probed a pretrained RoBERTa-large model. The same training policies from the fine-tuning step were followed in this stage. Two different classifiers were tested as probes:
  • 1-layer A single affine transformation is applied to the last hidden state of the first token (the [CLS] token). The classification head has shape (1024, labels). We used a dropout of 0.5 before the classifier.
@@
Both probes led to similar results. We take this as strong evidence that the knowledge used in the logical reasoning tasks can be linearly recovered from the internal representations of RoBERTa-large.
Table 3: Results for the cross-probing task: On the left, we present the best fine-tuned RoBERTa-large models for each dataset. The datasets used for training and testing the probes are listed on the top. We report only the best result for each probe. Blue cells indicate the probe of a fine-tuned model on the same dataset. In parentheses, we denote the percentage difference from the pretrained model. “Largest class” refers to the score for selecting the class with the highest frequency in the training set.
@@
Larg. class 35.29 25.00 50.02 50.03
Table 3 displays the best results attained in the cross-probing task. The left column lists the best RoBERTa-large fine-tuned models from the previous experiment, while the remaining columns represent the datasets they were probed against. The blue cells along the diagonal indicate instances where models were probed on the same datasets they were initially fine-tuned on. Percentage differences in accuracy relative to the pretrained case are reported in parentheses.
As expected, the fine-tuned model for a specific dataset performed better in that same dataset, albeit less than in the fine-tuning scenario. This makes sense, since in fine-tuning, both the model’s body and head are optimized, whereas in a probe the head is in charge of all the learning. These results serve as a sanity check that our probing is working.
The first row in Table 3 contains the scores for the probes with pretrained RoBERTa. Accuracy levels for LogicNLI and RuleTaker closely resemble the largest class baseline, while the result for FOLIO falls below its corresponding baseline. Pretrained RoBERTa only helped to solve SimpleLogic, a dataset constrained to conjunctive implication with minimal language variation. Its pretraining scheme, dynamic masked language modeling, did not equip it with adequate logic-like knowledge to solve more complex reasoning problems without specific training.
Clearly, if logical reasoning was not in pretrained RoBERTA, we would expect it to have been acquired in the fine-tuning process. After all, RoBERTa was able to solve the hypothesis classification tasks after training. When analyzing the results of the cross-probing task, however, we may have some doubt as to whether a general logical reasoning ability in fact emerged from fine-tuning.
In general, the fine-tuned LMs showed limited transferability when probed on different datasets. Although some gain was achieved as compared to the pretrained model, they remained well below what the LM fine-tuned on the same dataset obtained. SimpleLogic presents a singular case. The LMs fine-tuned on the other datasets performed similarly to the pretrained model in this dataset. This is despite SimpleLogic covering only a subset of propositional calculus, a domain included in those datasets. Conversely, the LM originally fine-tuned on SimpleLogic did not exhibit improved performance on other datasets, indicating a lack of acquired general logical reasoning ability during its training.
Two main conclusions can be taken from this experiment:
@@
    If the difference in accuracy between pretrained RoBERTa and the largest class baseline is indicative of the amount of logical reasoning contained in this LM, then pretrained RoBERTa’s logical reasoning abilities are very limited.
@@
    If the difference in accuracy between fine-tuned LMs (when applied to different datasets) and the pretrained RoBERTa model is indicative of the amount of logical reasoning they acquired in the fine-tuning process, then these LMs have acquired little or no general logical reasoning capability — suggesting that they mostly learned statistical features of the datasets. This is in line with other findings for transformer models (Zhang et al., 2022; Dziri et al., 2023).
@@ 6 Inspecting Fine-Tuned Models Layerwise
We now address another question: which parts of the LMs are more capable of solving logical reasoning tasks? To answer that, we probe the different layers of the fine-tuned RoBERTa-large models using the same datasets they were trained on. Our goal is to identify which layers are more effective in deducing hypotheses. Similar to the previous experiment, the fine-tuned LMs were frozen, and a classifier was trained on top of each layer. However, in this case, the fine-tuned LMs were only probed using the datasets they were previously trained on. The same classifiers from the last experiment were tested. They were positioned on top of the [CLS] token of each layer, with 25 layers in total (24 transformer blocks plus the initial embedding layer). We adopted the same configurations from the previous tests: models were trained with early stopping for up to 50 epochs (patience of 5 epochs) based on the validation loss, using two learning rates (1e-5 and 1e-6).
@@
Figure 2: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 1-layer classifier and a learning rate of 1e-6.
[…]
A similar behavior is exhibited by all models. They remain close to the pretrained baseline in the low and mid layers. Accuracy then grows rapidly in the final layers, achieving a performance equal to the cross-probing baseline.
@@
How does that compare to other types of knowledge found on transformer models? Literature on transformers show that surface information is mostly captured by lower layer information, e.g., sentence length (Jawahar et al., 2019); middle layers are responsible for processing syntactic knowledge, e.g., syntax trees (Hewitt and Manning, 2019); and higher layers are responsible for task-specific functions (Hao et al., 2019) and contextual representations (Ethayarajh, 2019).
@@
In our layerwise probing, higher layers were the only layers ones that were able to solve the hypothesis classification task better than a pretrained model. This suggests that the knowledge acquired in the fine-tuning training was connected to dataset-specific features, rather than general representations. It also explains why the information was not transferable among datasets.
SimpleLogic was the only case that presented a growth in the initial layers. This behavior may indicate that the dataset is solvable through the use of some heuristics based on shallow statistical features, such as the number of premises, as discussed by Zhang et al. (2022).
@@ 7 Conclusion
Logical reasoning is a valuable ability that humans use in thinking, arguing, and planing; it is also a core component of many AI systems. We investigated here the role of logical reasoning in encoder-only transformer models. By gathering a number of logical reasoning datasets, we saw that LMs can be trained to perform complex logical tasks with relative success. However, upon closer inspection, doubts arose regarding whether these models have truly learned to reason according to logical rules. First, by probing a pretrained RoBERTa-large model with logical reasoning datasets, it became apparent that this language model did not possess intrinsic logical reasoning abilities. Second, models fine-tuned on one dataset struggled to generalize well to others, even within the same domain. This observation suggests that these language models did not acquire robust logical reasoning capabilities. Third, the knowledge necessary to solve logical reasoning tasks seems to emerge primarily at higher, more contextual layers, which appear to be linked to statistical features of the datasets rather than deeper representations.
@@ 8 Limitations
We run experiments for a large variety of LMs in Section 4; however, due to space and time constraints, we focused on RoBERTa-large for our analysis in Sections 5-6. While we expect the same behavior to appear in other encode-only transformer models, further tests are needed to verify whether conclusions can be reliably extended to them. We have not explored decoder nor encoder-decoder models either, as those are better tested through behavioral tests, rather than structural probes. Furthermore, it could be the case that robust logical reasoning is an emergent ability only manifested in large language models (Wei et al., 2022).
In a different direction, other types of representations could be analyzed in relation to logical reasoning, such as attentions and feedforward computations. Further work should also concentrate on other types of logical formalism.
Appendix A Datasets
@@
(Han et al., 2022) is a human-annotated dataset for FOL reasoning problems. Logically-sound contexts were generated in two ways: in the first, annotators created contexts from scratch, based on random Wikipedia pages; in the second, a template of nested syllogisms was used, from which annotators then replaced the abstract entities and categories by nouns, phrases or clauses, as to make the text to reflect real-life scenarios. Next, the authors verified the alignment between natural language sentences and FOL formulas and added implicit commonsense knowledge as premises. After that, they verified the syntactic validity and label consistency of FOL formula annotations with a FOL prover. Finally, sentences were reviewed for grammar issues and language fluency. Only train and validation tests are available, so we used the latter for reporting tests. Hypotheses can be True, False, or Unknown.
@@
(Tian et al., 2021) is a FOL dataset created through a semi-automatic method. A set of logical templates was defined and then filled by subjects and predicates sampled from predefined sets. Next, manual edits were made to correct grammatical errors and resolve semantic ambiguities. Hypothesis are classified as Entailment, Contradiction, Neutral, and Paradox. A paradox is defined as a situation where both a sentence and its contrary can be inferred from the premises. We used the standard version of the dataset, which encompasses all labels.
@@
(Clark et al., 2020) is a logical reasoning dataset in which rules are conjunctive implications. Predicates may be negated and facts may be either attributes (which assign properties to entities) or relations (which relate two entities). We used the ParaRules version, where rules and facts were paraphrased by crowdworkers into more natural language; paraphrased constructions were then combined to form new templates. We also used the updated version of RuleTaker (problog), which eliminated some world model inconsistencies. Hypothesis can be True, when a hypothesis follows from the premises, and False otherwise (closed-world assumption, CWA).
@@
(Zhang et al., 2022) is similar to RuleTaker, only supporting conjunctive implication (facts are simply conjunctive implications with empty antecedents). Language variance is virtually removed by using a fixed template for translating FOL into natural language and by the use of a small random list of words as predicates. Argumentative complexity is limited by setting thresholds for input length, number of predicates, and reasoning depth. We reconstructed the original template based on the examples given in the paper. For our tests, we used the RP Balanced version. We undersampled the largest class (True) to obtain the same number of observations as for the False class. Labels can be True and False (CWA).
@@
LogicNLI https://github.com/omnilabNL
/LogicNLI
@@
SimpleLogic https://github.com/joshuacnf
/paradox-learning2reason
SNLI https://nlp.stanford.edu
/projects/snli/
MultiNLI https://cims.nyu.edu
/~sbowman/multinli/
SciTail https://huggingface.co
/datasets/scitail
@@
Appendix B Label distribution
@@
Appendix C Laywerwise probing
@@
Figure 3: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 3-layer classifier and a learning rate of 1e-6.
Figure 4: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 1-layer classifier and a learning rate of 1e-5.
Figure 5: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 3-layer classifier and a learning rate of 1e-5.

### v2

@@ 5 Cross-Probing Fine-Tuned LMs
The encoder-only transformer models showed reasonable performance on the hypothesis classification task, where they were fine-tuned on the logical reasoning datasets. This, however, raises some questions: has the ability to solve this task, whatever it is, been acquired during the fine-tuning stage, or was it present from the start (i.e., from pretraining)? Most importantly, have LMs truly developed a generalized logical reasoning capability? To examine these questions, we run a cross-probing task: we take the LMs previously fine-tuned on our logical datasets, as well as a pretrained LM, and probe them on these same datasets. Given the large number of possible tests, we restricted our investigation to a single model, RoBERTa-large, so as to dig deeper on it. This LM demonstrated a suitable balance between performance, consistency among datasets, and training time in the previous tests.
To start, we took the best fine-tuned RoBERTa-large model for each dataset and removed their classification heads, leaving just the transformer blocks, as in a pretrained model. Then, we attached a new classification head to it; i.e., the probe. As in the fine-tuning stage, we passed the formatted inputs in natural language to the LMs and tried to predict the correct label for a set of premises. However, unlike the fine-tuning stage, only the probe is updated now, while the model’s body is kept frozen during the backward pass. The goal is to assess if some logical reasoning ability was learned by the LM without letting the model adapt to the task.
The same training policies from the fine-tuning step were followed in this stage: models were trained with early stopping for up to 50 epochs (patience of 5 epochs) based on the validation loss, using two learning rates (1e-5 and 1e-6). Also, two different classifiers were tested as probes:
  • 1-layer A single affine transformation is applied to the embedding of the [CLS] token. The classification head has shape (1024, labels); 1024 being the dimensionality of RoBERTa-large hidden states. We used a dropout of 0.5 before the classifier.
@@
In the tests, the two probes led to similar results. We take this as strong evidence that the knowledge used in the logical reasoning tasks, whatever it is, can be linearly recovered from the internal representations of RoBERTa-large.
Table 3: Results for the cross-probing task. On the left, we present the best fine-tuned RoBERTa-large models for each dataset. The datasets used in the probes are listed at the top. We report only the best result for each probe. Blue cells indicate the probe of a fine-tuned model on the same dataset. In parentheses, we denote the percentage difference from the pretrained model. “Largest class” refers to the accuracy achieved by always selecting the class with the highest frequency in the training set.
@@
Largest class 35.29 25.00 50.02 50.03
Table 3 displays the best results attained in the cross-probing task. The left column lists the RoBERTa-large fine-tuned models from the previous experiment, while the remaining columns represent the datasets they were probed against. The blue cells along the diagonal indicate instances where models were probed on the same datasets they were initially fine-tuned on. Percentage differences in accuracy relative to the pretrained case are reported in parentheses. As expected, the fine-tuned model for a specific dataset performed better in that same dataset, albeit less than in the fine-tuning scenario. This makes sense, since in fine-tuning, both the model’s body and head are optimized, whereas in a probe the head is in charge of all the learning. These results serve as a sanity check that our probing is working.
The first row in Table 3 contains the scores for the probes with pretrained RoBERTa. Accuracy levels for LogicNLI and RuleTaker closely resemble the largest class baseline, while the result for FOLIO falls below its corresponding baseline. Pretrained RoBERTa only helped to solve SimpleLogic, a dataset constrained to conjunctive implication with minimal language variation. Its pretraining scheme, dynamic masked language modeling, did not equip it with adequate logic-like knowledge to solve complex reasoning problems without specific training.
We can see from this that pretrained RoBERTa appears to have no proper logical reasoning skills. But has it acquired such an ability through fine-tuning on logical datasets? After all, RoBERTa-large was able to solve the hypothesis classification tasks reasonably well after specific training. When analyzing the results of the cross-probing task, however, we may doubt whether a general logical reasoning ability in fact emerged from fine-tuning.
In general, the fine-tuned LMs showed limited transferability when probed on different datasets. Although some gain was achieved compared to the pretrained model, they remained well below what an LM fine-tuned on the same dataset could obtain. SimpleLogic presents an interesting case. The LMs fine-tuned on the other datasets performed similarly to the pretrained model on this dataset. This is despite SimpleLogic covering only a subset of propositional calculus, a domain included in those datasets. One would expect that a model capable of solving more complex problems would be able to reason on this simpler dataset (in terms of logical scope and linguistic variability). At the same time, the LM fine-tuned on SimpleLogic did not exhibit improved performance on the other datasets, indicating a lack of acquired general logical reasoning ability during its training.
Two main conclusions can be drawn from this experiment:
@@
    If the difference in accuracy between pretrained RoBERTa and the largest class baseline indicates the amount of logical reasoning contained in this LM, then pretrained RoBERTa seems to have very limited or no logical reasoning abilities.
@@
    If the difference in accuracy between fine-tuned LMs (when applied to different datasets) and the pretrained RoBERTa model indicates the amount of logical reasoning they acquired in the fine-tuning process, then these LMs have acquired little or no general logical reasoning capability as well, suggesting that they mostly learned statistical features of the datasets. This aligns with other findings for transformer models [55, 12].
@@ 6 Inspecting Fine-Tuned Models Layerwise
We now address another question: which parts of the LMs are more capable of solving logical reasoning tasks? To answer this, we probe the different layers of the fine-tuned RoBERTa-large models using the same datasets they were trained on. Our goal is to identify which layers are more effective in deducing hypotheses. We expect this to provide further evidence of what sort of knowledge LMs are using to solve the hypothesis classification task. Similar to the previous experiment, the fine-tuned LMs were frozen, the older classification head was removed, and a probe was trained on top of the layers. More concretely, for each layer i in the model, we passed the premises through the model up to layer i, and used the outputted embedding of the [CLS] token at that layer as the input to the probe. As in the previous experiment, only the probe was trained. The same two classifiers from the last experiment were tested. They were positioned on top of the [CLS] token of each layer, with 25 layers in total (24 transformer blocks plus the initial embedding layer). We adopted the same configurations from the previous tests: models were trained with early stopping for up to 50 epochs (with a patience of 5 epochs) based on the validation loss, using two learning rates (1e-5 and 1e-6).
@@
Figure 2: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by gray lines, while the values achieved in the cross-probing task are represented by black lines. The colored bars indicate the change in accuracy from the previous layers. Probing was performed with a 1-layer classifier and a learning rate of 1e-6.
Figure 2 shows the accuracy for the probes stacked on the various layers of the fine-tuned LMs, using a 1-layer classifier and a learning rate of 1e-6 (Appendix 0.C presents graphs for the other configurations). The blue line plots the accuracy on the task for each layer, while the bars display the differential score for layeri; i.e., the change in accuracy from layeri to layeri-1 [45]. The gray line marks the pretrained model baseline, and the black line indicates the score achieved in the cross-probing task.
@@
A similar behavior is exhibited by all models. They remain close to the pretrained baseline in the low and mid layers. Accuracy then grows rapidly in the final layers, achieving a performance equal to the cross-probing baseline. How does this compare to other types of knowledge found in transformer models? Literature on transformers shows that surface information, such as sentence length [23], is mostly captured by lower layers. Middle layers are responsible for processing syntactic knowledge, like syntax trees [21]. Finally, higher layers are responsible for task-specific functions [18] and contextual representations [13].
@@
In our layerwise probing, higher layers were the only ones able to solve the hypothesis classification task better than a pretrained model. This suggests that the knowledge acquired during fine-tuning was connected to dataset-specific features rather than general representations. It also explains why the information was not transferable among datasets. Although indirect, this experiment provides further evidence that encoder-only transformer models do not possess robust logical reasoning capabilities. SimpleLogic was the only case that presented a growth in the initial layers. This behavior may indicate that the dataset is solvable through the use of some heuristics based on shallow statistical features, such as the number of premises, as discussed by [55].
@@ 7 Conclusion
Logical reasoning is a valuable ability that humans use in thinking, arguing, and planning, as well as a core component of many AI systems. Here, we investigated the role of logical reasoning in encoder-only transformer models. By gathering a number of logical reasoning datasets, we observed that language models can be trained to perform complex logical tasks with relative success. However, upon closer inspection, doubts arose regarding whether these models have truly learned to reason according to logical rules. First, by probing a pretrained RoBERTa-large model with logical reasoning datasets, it became apparent that this language model did not possess intrinsic logical reasoning abilities. Second, models fine-tuned on one dataset struggled to generalize well to other datasets, even within the same domain. This observation suggests that these language models did not acquire robust logical reasoning capabilities even after specific training. Third, the knowledge necessary to solve logical reasoning tasks seems to emerge primarily at higher, more contextual layers, probably linked to statistical features of the datasets rather than deeper representations.
@@ 8 Limitations
We run experiments for a large variety of encoder-only transformer models in Section 4. However, due to space and time constraints, we focused on RoBERTa-large for the analysis in Sections 5-6. While we expect the same behavior to appear in the other encode-only models, further tests are needed to verify whether conclusions can be reliably extended to them. We have not explored decoder nor encoder-decoder models either, which could widely extend the number of models to be tested. We cannot rule out the possibility that robust logical reasoning is an emergent ability only manifested in large language models [50]. Additionally, other types of representations, such as attentions and feedforward computations, could be analyzed in relation to logical reasoning. Further work should also focus on other types of logical formalism beyond PC and FOL.
Acknowledgements
This work was supported by the Center for Artificial Intelligence USP/IBM/FAPESP (C4AI), jointly funded by the São Paulo Research Foundation (FAPESP grant 2019/07665-4) and by the IBM Corporation. Research by Marcos José has been carried out with support by Itaú Unibanco S.A. through the scholarship program Programa de Bolsas Itaú (PBI) ; Fabio Cozman was partially supported by CNPq grants 312180/2018-7 and 305753/2022-3. We acknowledge support also by CAPES - Finance Code 001.
Appendix 0.A Datasets
@@
[17] is a human-annotated dataset for FOL reasoning problems. Logically-sound contexts were generated in two ways: in the first, annotators created contexts from scratch, based on random Wikipedia pages; in the second, a template of nested syllogisms was used, from which annotators then replaced the abstract entities and categories by nouns, phrases or clauses, as to make the text to reflect real-life scenarios. Next, the authors verified the alignment between natural language sentences and FOL formulas and added implicit commonsense knowledge as premises. After that, they verified the syntactic validity and label consistency of FOL formula annotations with a FOL prover. Finally, sentences were reviewed for grammar issues and language fluency. Only train and validation tests are available, so we used the latter for reporting tests. Hypotheses can be True, False, or Unknown.
@@
[47] is a FOL dataset created through a semi-automatic method. A set of logical templates was defined and then filled by subjects and predicates sampled from predefined sets. Next, manual edits were made to correct grammatical errors and resolve semantic ambiguities. Hypothesis are classified as Entailment, Contradiction, Neutral, and Paradox. A paradox is defined as a situation where both a sentence and its contrary can be inferred from the premises. We used the standard version of the dataset, which encompasses all labels.
@@
[4] is a logical reasoning dataset in which rules are conjunctive implications. Predicates may be negated and facts may be either attributes (which assign properties to entities) or relations (which relate two entities). We used the ParaRules version, where rules and facts were paraphrased by crowdworkers into more natural language; paraphrased constructions were then combined to form new templates. We also used the updated version of RuleTaker (problog), which eliminated some world model inconsistencies. Hypothesis can be True, when a hypothesis follows from the premises, and False otherwise (closed-world assumption, CWA).
@@
[55] is similar to RuleTaker, only supporting conjunctive implication (facts are simply conjunctive implications with empty antecedents). Language variance is virtually removed by using a fixed template for translating FOL into natural language and by the use of a small random list of words as predicates. Argumentative complexity is limited by setting thresholds for input length, number of predicates, and reasoning depth. We reconstructed the original template based on the examples given in the paper. For our tests, we used the RP Balanced version. We undersampled the largest class (True) to obtain the same number of observations as for the False class. Labels can be True and False (CWA).
@@
LogicNLI https://github.com/omnilabNL/LogicNLI
@@
SimpleLogic https://github.com/joshuacnf/paradox-learning2reason
@@
Appendix 0.B Label distribution
@@
Appendix 0.C Laywerwise probing
@@
Figure 3: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by gray lines, while the values achieved in the cross-probing task are represented by black lines. The colored bars indicate the change in accuracy from the previous layers. Probing was performed with a 3-layer classifier and a learning rate of 1e-6.
Figure 4: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by gray lines, while the values achieved in the cross-probing task are represented by black lines. The colored bars indicate the change in accuracy from the previous layers. Probing was performed with a 1-layer classifier and a learning rate of 1e-5.
Figure 5: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by gray lines, while the values achieved in the cross-probing task are represented by black lines. The colored bars indicate the change in accuracy from the previous layers. Probing was performed with a 3-layer classifier and a learning rate of 1e-5.